# Subword models struggle with word learning, but surprisal hides it

### v1

@@ Abstract
We study word learning in subword and character language models with the psycholinguistic lexical decision task. While subword LMs struggle to discern words and non-words with high accuracy, character LMs solve this task easily and consistently. Furthermore, when comparing word learning and syntactic learning, both processes are separable in character LM where word learning predates syntactic learning, whereas these processes are simultaneous in subword LM. This raises questions about the adequacy of subword LMs for modeling language acquisition and positions character LMs as a viable alternative.
@@ 1 Introduction
When humans acquire first language(s), they first learn to recognize single words before understanding the grammatical processes governing them Tomasello (1992); Behrens (2021). This simple fact about language acquisition has found surprisingly little attention in the increasing amount of work that treats LMs as models of language learners Warstadt and Bowman (2022); Portelance and Jasbi (2024). While word learning in children is well studied, the implicit word learning processes in LMs are not. Current studies overwhelmingly focus on syntax Mueller et al. (2022); Choshen et al. (2022), or investigate word learning in close connection to syntax through surprisal Chang and Bergen (2022); Portelance et al. (2023); Shafiabadi and Wisniewski (2025); Ficarra et al. (2025). Architecture-wise, a key limitation to the precise study of word learning is subword tokenization (e.g. BPE, Gage, 1994), which splits words into linguistically and cognitively implausible units Beinborn and Pinter (2023); Arnett and Bergen (2025).
Figure 1: Word learning in natural/artificial learners
To gauge word learning in a syntax-independent manner, we use the psycholinguistic lexical decision task Meyer and Schvaneveldt (1971); Le Godais et al. (2017), i.e. deciding which word in given word/non-word pair is real (e.g. doggie vs. moggie, cf. Figure 1). We find that models with character-level tokenization and therefore more fine-grained representations learn this task quickly and reliably. In contrast, subword LMs of all sizes perform much worse in a syntax-independent lexical decision setting and only reach comparable accuracy when stimuli are measured through surprisal, “unexpectedness” in syntactic contexts. By comparing word and syntactic learning (measured via BLiMP, Warstadt et al., 2020), we further find that character models quickly acquire lexical knowledge and only later develop syntactic knowledge. In subword models, however, word learning happens later and concurrently with syntax learning, bringing further evidence against the cognitive plausibility of subword tokenization. This shows how elementary decisions (like choice of tokenization) can tremendously influence the learning dynamics and trajectories that can be observed in LMs, a fact that should receive more scrutiny in studies of LMs as models of language acquisition.
@@ 2 Related work
Word learning in LMs is most commonly measured via word-level surprisal (negative log-probability, Hale, 2001). Chang and Bergen (2022) define a surprisal threshold below which words are said to be learned. They find that BERT and GPT-2 learn frequent function words early, unlike children, who first utter nouns and verbs. Portelance et al. (2023) show that in LSTMs trained on child-directed speech, surprisal correlates with word-level age of acquisition. Chang et al. (2024) observe that learning curves for surprisal values are stable for frequent tokens, while infrequent tokens are “forgotten” again over pre-training. Shafiabadi and Wisniewski (2025) introduce anti-surprisal (in incorrect contexts) to track false usage, which also fluctuates over pre-training. These studies cast word learning as the ability to anticipate words given a preceding sequence of tokens, i.e. estimating its expectedness in a given syntactic (and semantic) context. We note a certain conceptual leap to the original works on surprisal, where it is primarily viewed as an incremental measure of processing difficulty in syntactic comprehension Levy (2008); Demberg and Keller (2009). A simple word like dog might be surprising and therefore hard to parse in some contexts, but very expected in others, independently of it being already learned on the pure word level. A further methodological drawback of surprisal as a measure of word learning is that it corresponds almost directly to the next-token prediction objective LMs are trained on. This contrasts with typical probing paradigms used, e.g., in the domain of syntax, which often implement the idea to “challenge” models in minimal pair-style set-ups that cannot be observed directly in training, thereby testing implicit linguistic knowledge rather than observed patterns in corpus data. In a similar vein, we want to probe the word knowledge of an LM at a fundamental level, and beyond surface-level word sequence patterns that LMs are known to excel in predicting. We want to know if the learner knows that the word doggie exists in the English language, but moggie does not.
@@
While lexical decision has been used to study lexical abilities in humans, it remains an underexplored benchmark. Le Godais et al. (2017) show that character-based LSTMs achieve about 95% accuracy on such tasks. Lavechin et al. (2023) find that speech-based LMs need significantly more input to still perform poorly (56.8%) than phoneme-level LSTMs (75.4%) on a phonetic dataset. For the same data, Goriely et al. (2024) find that regular GPT-2 models achieve 70% accuracy and a character-based GPT-2 reaches nearly 90%. Bunzeck et al. (2025) report near-perfect lexical decision accuracy for regular Llama models, while phoneme models perform at 60–70%. None of these studies report learning trajectories.
3 Models, data, and pre-training
We train triplets of increasingly larger Llama models Touvron et al. (2023) with character/subword tokenization on the BabyLM 10M corpus Choshen et al. (2024), which contains developmentally plausible data. Hyperparameters, losses, etc., are found in Appendix A. As additional ablations, we test subword Pythias Biderman et al. (2023) and character/subword GPT-2 models (Goriely et al., 2024).
@@ 3 Models, data, and pre-training
Subword (BPE) Pythia 14M 825GB 66.6% 62.5% 90.5% 85.5% 71.4% 77.7%
70M 72.5% 68.8% 94.5% 94.0% 77.0% 83.6%
160M 77.8% 73.0% 96.4% 95.8% 78.0% 85.7%
410M 81.9% 78.1% 97.7% 97.9% 77.1% 84.1%
1B 87.5% 83.2% 97.7% 97.9% 76.6% 83.8%
1.4B 87.8% 81.6% 97.9% 97.9% 76.5% 84.7%
GPT-2 97.5M 100M words 35.6% 79.1% 99.0% 99.2% 84.7% 86.9%
Llama 2.51M 10M words 70.9% 58.4% 86.7% 70.9% 78.6% 67.7%
7.77M 79.5% 63.2% 91.3% 78.1% 81.1% 72.9%
30.03M 83.6% 68.6% 92.7% 81.1% 83.7% 76.1%
Character GPT-2 85.3M 100M words 98.7% 97.3% 99.8% 99.4% 98.0% 96.3%
Llama 0.49M 10M words 97.6% 83.0% 98.2% 84.3% 98.0% 83.1%
3.73M 98.9% 90.2% 99.4% 90.3% 98.5% 88.8%
21.94M 99.0% 93.3% 99.8% 94.7% 99.0% 92.5%
Table 1: Accuracy for lexical decision (i) without context, (ii) measured through surprisal and (iii) anti-surprisal
@@ 4 Experiments
We follow the idea of forced-choice lexical decision Baddeley et al. (1993); Le Godais et al. (2017), where participants are presented with an existing word and a synthesized non-word, and must decide which is real. To test LMs at this, we generate minimal pairs of words/non-words, akin to established linguistic benchmarks such as BLiMP, with wuggy Keuleers and Brysbaert (2010). We derive 1,000 non-words (e.g. monding) each from 1,000 high-frequency/low-frequency words (e.g. sending), which preserve syllable-bigram frequencies, match their origin words in length and look graphematically plausible (more details in Appendix B).
@@
To measure LMs’ knowledge of words presented in regular syntactic contexts, we calculate the surprisal of words and non-words $(w,*w)$ as $-log(P(w_{i}|w_{n<i}))$, i.e. the degree to which the LM is “surprised” by the word in the context of plausible preceding tokens. We create stimuli by sampling sentences that contain the existing words from OpenSubtitles (Lison and Tiedemann, 2016) and substituting them with matching non-words for the false stimuli. If $-log(P(w_{i}|w_{n<i}))<-log(P(*w_{i}|w_{n<i}))$, the LM’s decision is correct.
@@
We create negative samples by selecting sentences that our original words do not occur in, and then randomly placing words/non-words into these sentences at the same index. By doing so, we compromise between lexical decision and surprisal measurement. Words are presented in the context of existing words, but without plausible syntactic/semantic signals in the surrounding data. In line with Shafiabadi and Wisniewski (2025), we call this measure anti-surprisal. In contrast to them (and Ficarra et al., 2025), however, we are not interested in whether models are surprised by unfitting words in context, but only want to provide some positive lexical signal (without other signals). Again, if $-log(P(w_{i}|w_{n<i}))<-log(P(*w_{i}|w_{n<i}))$, the LM’s decision is correct.
@@
To assess when word learning happens in relation to syntax learning, we further evaluate intermediate checkpoints of our models on our lexical tests and BLiMP as a syntactic benchmark. In line with previous studies Chang and Bergen (2022); Viering and Loog (2023), we space our checkpoints logarithmically – 10 for the first 10% of training, 9 more for the remaining 90%. For the Pythia models, we extract similarly spaced checkpoints (the GPT-2 models do not provide checkpoints, so we exclude them).
@@ 5 Results
The lexical decision results (Table 1) show a strong contrast between character and subword models. Character models achieve near-perfect accuracy (97–99%) on high-frequency words, regardless of model size. The performance on low-frequency words steadily increases with model size and reaches a near-perfect level for our largest character Llama and the character GPT-2. On the other hand, all BPE models get surprisingly low scores on high-frequency words: The smallest Pythia model discriminates between word and non-words with an accuracy of only 67%, and the BPE GPT-2 actually performs below the chance baseline. Even the largest and best BPE model reaches only 87.8% on high-frequency words – almost 10% less than the smallest character model.
Scaling laws generally hold, with larger models outperforming smaller ones. Interestingly, in the BPE models there is a consistent gap between high and low-frequency words which cannot be closed by the larger model sizes. The smaller character models also show a significant gap between high and low-frequency words, but this gap narrows considerably with larger model sizes. These results point to substantial differences in how BPE and characters models learn words. Such a surprising lack of ability in distinguishing words from non-words (where no context is given) is a blatant, hitherto overlooked gap in BPE-based LMs.
@@
Results for the second experiment (Table 1) differ from those for lexical decision. In the surprisal setting, the difference between BPE and character models is less pronounced. On high-frequency data, nearly all models (except the smallest BPE Llama) achieve over 90% accuracy. Still, larger character models yield the very best results. In the low-frequency data condition, the pattern is similar, though scores are generally lower. Very large BPE models outperform our Llamas there, but the character GPT-2 remains superior. This may be attributed to the limited lexical exposure of our Llamas, trained on only 10M tokens. In the anti-surprisal setting, character models again drastically outperform BPE models and achieve nearly perfect scores on high-frequency data, while BPE models only reach 70–80% accuracy. This setting is the only one where Pythia models get better scores on low-frequency data, with a gap of 6–8%, which increases with model size (not the case for BPE Llamas or character models). This is another indicator for the entanglement of word learning and syntactic learning in subword models. It is plausible that, for high-frequency words, the BPE models have strong expectations about which word should come next in a certain context, and because this expectation is not matched by the real (but ill-fitting) word, a made-up none word is generally preferred. In any case BPE models require context to perform lexical decision tasks consistently, catching up to character models if (and only if) provided with additional syntactic/semantic information. While random context somewhat aids BPE models, a substantial gap remains between the largest BPE models and the character models, where performance remains excellent, even in the presence of implausible contexts.
@@
We display learning curves for syntactic agreement phenomena, lexical decision and both surprisal conditions in Figure 2 (complete curves reproduced in Appendix C). For character models, the high-frequency, low-frequency, and syntactic curves are clearly separated. On high-frequency data, word learning is rapid and follows power-law curves; the low-frequency scores improve a little later and at a lower rate, but with the same trajectories. Syntactic phenomena improve later, mostly in s-shaped curves (e.g. det.-noun agreement), or are not learned at all in small models (subj.-verb agreement). In contrast, the syntactic and lexical curves for BPE models form sheaves of s-shaped trajectories. There is no principle difference in learning dynamics between the syntactic and the lexical level, [^2] To corroborate these findings, we also calculate Spearman-rank correlation between ordered accuracies for lexical tasks and BLiMP paradigms (Results in Appendix D). improvements occur simultaneously. This further confirms the results of our previous experiments: in BPE models, word learning is extremely dependent on syntax learning and words cannot be recognized reliably outside of plausible contexts. Additionally, these different levels of learning cannot be disentangled in BPE models, whereas in character models syntactic learning follows word learning. [^3] Interestingly, the learning process is not finished when accuracy curves stabilize. We provide an analysis of the absolute differences between response values in Appendix F.
6 Discussion and conclusion

### v2

@@ Abstract
We study word learning in subword and character language models with the psycholinguistic lexical decision task. While subword LMs struggle to discern words and non-words with high accuracy, character LMs solve this task easily and consistently. Only when supplied with further contexts do subword LMs perform similarly to character models. Additionally, when looking at word-level and syntactic learning trajectories, we find that both processes are separable in character LMs. Word learning happens before syntactic learning, whereas both occur simultaneously in subword LMs. This raises questions about the adequacy of subword LMs for modeling language acquisition and positions character LMs as a viable alternative to study processes below the syntactic level.
@@ 1 Introduction
When humans acquire their first language(s), they first learn to recognize single words, mostly from short, fragmentary utterances Cameron-Faulkner et al. (2003); Bunzeck and Diessel (2024), before fully understanding the grammatical processes governing them Tomasello (1992); Behrens (2021). This simple fact about language acquisition has received surprisingly little attention in the body of work that treats LMs as models of language learners Warstadt and Bowman (2022); Portelance and Jasbi (2024). While word learning in children is comparatively well studied (Plunkett, 1997; Yu and Ballard, 2007; Waxman and Gelman, 2009; Bergelson and Swingley, 2012; Clark and Casillas, 2015; Frank et al., 2021), the implicit word learning processes in LMs are not. Current studies focus on syntax Mueller et al. (2022); Choshen et al. (2022), or investigate word learning in close connection with syntax through surprisal Chang and Bergen (2022); Portelance et al. (2023); Shafiabadi and Wisniewski (2025); Ficarra et al. (2025). Architecture-wise, a key limitation to the precise study of word learning is subword tokenization (e.g., BPE, Gage, 1994), which splits words into linguistically Arnett and Bergen (2025) and cognitively implausible units Beinborn and Pinter (2023).
Figure 1: Illustration of word learning in human learners and transformer LLMs (top), and of our lexical decision test that probes discrimination of words from non-words (bottom). While human learners build up an mental lexicon from experience with language, artificial learners assign probabilities to strings based on their frequency.
To gauge word learning in a syntax-independent manner, we use the psycholinguistic lexical decision task Meyer and Schvaneveldt (1971); Le Godais et al. (2017), i.e., deciding which word in a given word/non-word pair is real. We find that models with character-level tokenization learn this task quickly and reliably. In contrast, subword LMs of all sizes perform significantly worse in a syntax-independent setting and only achieve comparable accuracy when stimuli are measured through surprisal, or “unexpectedness,” in linguistic context. By comparing word and syntactic learning (measured via BLiMP, Warstadt et al., 2020), we further find that character models quickly acquire word knowledge and only later develop syntactic knowledge. In subword models, word and syntax learning happen concurrently. This demonstrates how elementary modeling decisions, such as tokenization methods, significantly impact learning trajectories in LMs, a fact that warrants more scrutiny when using LMs as models of language acquisition.
@@ 2 Related work
Word learning in humans is a multifaceted phenomenon that involves different kinds of (extra)linguistic knowledge Waxman and Gelman (2009): while phoneticians are concerned with word recognition and sequence segmentation Jusczyk (1999), developmental psychologists frequently equate word learning with correct reference to real world objects Stager and Werker (1997); Ackermann et al. (2020). Psycholinguists usually focus on the mental lexicon and learning which words belong to it Goldinger (1996), whereas usage-based scholars take into account children’s productions and their ability to be competent language users, even with few words Tomasello (1992).
@@
Although aspects like word recognition and sequence segmentation have been studied in LMs (e.g. Goriely and Buttery, 2025a), the most common approach to word learning in LMs is measuring the predictability of words via surprisal (negative log-probability, Hale, 2001). Chang and Bergen (2022) train LMs on book texts and wiki data. They define a surprisal threshold below which words are said to be learned and find that frequent function words are learned earliest. Here, the alignment between models and real learners is questionable: children first utter nouns and verbs Tomasello (2000), but also rely on function words for challenges like speech segmentation Dye et al. (2019). Portelance et al. (2023) show that in LSTMs trained on child-directed speech, surprisal correlates with word-level age of acquisition. Chang et al. (2024) observe that learning curves for surprisal values are stable for frequent tokens, while infrequent tokens are “forgotten” again over pre-training. Shafiabadi and Wisniewski (2025) introduce anti-surprisal (in incorrect contexts) to track false usage, which also fluctuates over pre-training. These studies cast word learning as the ability to anticipate words’ expectedness in a given syntactic (and semantic) context. We note a certain conceptual leap to the original works on surprisal, where it is primarily viewed as an incremental measure of processing difficulty in syntactic comprehension Levy (2008); Demberg and Keller (2009). A simple word like dog might be surprising and therefore hard to parse in some contexts, but very expected in others, independently of being already learned on the pure word level. A further methodological drawback of surprisal as a measure of word learning is that it corresponds almost directly to the next-token prediction objective LMs are trained on. This contrasts with typical probing paradigms used in the domain of syntax, which implement the idea to “challenge” models in minimal pair set-ups that are not observed directly as string sequences in training, thereby testing abstracted, implicit linguistic knowledge rather than observed patterns in the data. In a similar vein, we want to probe the word knowledge of an LM at a fundamental level and beyond surface-level word sequences that LMs are known to excel in predicting. We want to know if the artificial learner knows that the word doggie exists in the English language, but moggie does not.
@@
Lexical decision is widely used in human studies but remains an underexplored LM benchmark. Le Godais et al. (2017) show that character-based LSTMs achieve about 95% accuracy on such tasks. Lavechin et al. (2023) find that speech-based LMs need significantly more input to still perform poorly (56.8%) than phoneme-level LSTMs (75.4%) on a phonetic dataset. For the same data, Goriely et al. (2024) find that GPT-2-based subword BabyLMs achieve 70% accuracy and a comparable character models reach nearly 90%. Finally, for another lexical decision dataset, Bunzeck et al. (2025) report near-perfect accuracy for character-based grapheme Llama models, while phoneme models perform at 60–70%.
@@
Subword (BPE) Pythia 14M 825GB 66.6 62.5 90.5 85.5 71.4 77.7
70M 72.5 68.8 94.5 94.0 77.0 83.6
160M 77.8 73.0 96.4 95.8 78.0 85.7
410M 81.9 78.1 97.7 97.9 77.1 84.1
1B 87.5 83.2 97.7 97.9 76.6 83.8
1.4B 87.8 81.6 97.9 97.9 76.5 84.7
GPT-2 97.5M 100M words 35.6 79.1 99.0 99.2 84.7 86.9
Llama 2.51M 10M words 70.9 58.4 86.7 70.9 78.6 67.7
7.77M 79.5 63.2 91.3 78.1 81.1 72.9
30.03M 83.6 68.6 92.7 81.1 83.7 76.1
Character GPT-2 85.3M 100M words 98.7 97.3 99.8 99.4 98.0 96.3
Llama 0.49M 10M words 97.6 83.0 98.2 84.3 98.0 83.1
3.73M 98.9 90.2 99.4 90.3 98.5 88.8
21.94M 99.0 93.3 99.8 94.7 99.0 92.5
Table 1: Accuracy scores (in %) for (i) lexical decision, (ii) surprisal and (iii) anti-surprisal experiments
@@ 3 Experiments
Models
We train triplets of increasingly larger Llama models Touvron et al. (2023) with character/subword tokenization on the BabyLM 10M corpus Choshen et al. (2024). Training details are found in Appendix A. As ablations, we test subword Pythias Biderman et al. (2023) and character/subword GPT-2 models (Goriely et al., 2024).
Test data
We follow the idea of forced-choice lexical decision Baddeley et al. (1993), where participants must decide which is real: an existing word or a synthesized non-word. We use wuggy Keuleers and Brysbaert (2010) to generate minimal pairs of words/non-words that differ in one or two syllables, akin to syntactic minimal pair tests such as BLiMP. We derive 1,000 non-words (e.g. monding) each from 1,000 high-frequency/low-frequency words (e.g. sending), which preserve syllable-bigram frequencies and match their origin words in length (cf. Appendix B).
@@
To measure LMs’ knowledge of words presented in regular syntactic contexts, we calculate the surprisal of words and non-words $(w,*w)$ as $-log(P(w_{i}|w_{n<i}))$, i.e. the degree to which the LM is “surprised” by the word in the context of plausible preceding tokens, including a BOS token. We create stimuli by sampling sentences that contain our target words from OpenSubtitles (Lison and Tiedemann, 2016) and substituting them with matching non-words for the false stimuli. If $-log(P(w_{i}|w_{n<i}))<-log(P(*w_{i}|w_{n<i}))$, the LM’s decision is correct.
@@
Inspired by Shafiabadi and Wisniewski (2025), we include anti-surprisal, a measure of word surprisal on negative instances. We create negative samples by selecting sentences that our original words do not occur in, and then randomly [^1] In line with Shafiabadi and Wisniewski (2025), we do not further characterize the resulting violations on a syntactic or semantic level. placing words/non-words into these sentences at the same index, where $index\geq 3$. By doing so, we compromise between lexical decision and surprisal measurement. There are two reasons to include this measure: i) surprisal in negative samples provides the model with word material as context, but without semantic or syntactic signals that could prime the model towards recognizing it; this allows us to assess whether the mere presence of other words in context makes it easier for the model to distinguish words from non-words, compared to our lexical decision set-up where only whitespace is given. In addition, ii) we want to see if the presence of an ill-fitting context actively deteriorates performance in the sense of a model suddenly preferring non-words over existing words. Again, if $-log(P(w_{i}|w_{n<i}))<-log(P(*w_{i}|w_{n<i}))$, the LM’s decision is correct.
@@
To assess when word learning happens in relation to syntax learning, we further evaluate intermediate checkpoints of our models on our word learning tests and BLiMP as a syntactic benchmark. In line with previous studies Chang and Bergen (2022); Viering and Loog (2023), we space our checkpoints logarithmically – 10 for the first 10% of training, 9 more for the remaining 90%. For the Pythia models, we extract similarly spaced checkpoints (the GPT-2 models do not provide checkpoints, so we exclude them).
@@ 4 Results
The lexical decision results (Table 1) show a strong contrast between character and subword models. Character models achieve near-perfect accuracy (97–99%) on high-frequency words, regardless of model size. The performance on low-frequency words steadily increases with model size and reaches a near-perfect level for our largest character Llama and the character GPT-2. On the other hand, all BPE models get surprisingly low scores on high-frequency words: The smallest Pythia model discriminates between word and non-words with an accuracy of only 67%, and the BPE GPT-2 performs below the chance baseline. Even the largest and best BPE model reaches only 87.8% on high-frequency words – almost 10% less than the smallest character model.
Scaling laws generally hold, with larger models outperforming smaller ones. Interestingly, for the BPE models, there is a consistent gap between high and low-frequency words that cannot be closed by larger models. Smaller character models also show a performance gap between high and low-frequency words, but it narrows considerably with larger models. These results point to substantial differences in how subword and character models learn words. Such a surprising lack of ability in distinguishing words from non-words (without context) is a blatant, hitherto overlooked gap in subword models.
@@
Results for the second experiment (Table 1) differ from those for lexical decision. In the surprisal setting, the difference between BPE and character models is less pronounced. On high-frequency data, nearly all models (except the smallest BPE Llama) achieve over 90% accuracy. Still, larger character models yield the best results. In the low-frequency data condition, the pattern is similar, though scores are generally lower. Very large BPE models outperform our Llamas there, but the character GPT-2 remains superior. This may be attributed to the limited lexical exposure of our Llamas, trained on only 10M tokens. In the anti-surprisal setting, character models again drastically outperform BPE models and achieve nearly perfect scores on high-frequency data, while BPE models only reach 70–80% accuracy. This setting is the only one where Pythia models get better scores on low-frequency data, with a gap of 6–8%, which increases with model size (not the case for BPE Llamas or character models). This contrast between surprisal and antisuprisal might be an indicator of the entanglement of word learning and syntactic learning in subword models. It is plausible that for high-frequency words, the BPE models have strong expectations about which word should come next in a certain context, and because this expectation is not matched by the real (but ill-fitting) word, a made-up non-word is preferred. We argue that this should still not be the case for an ideal language model – if a model is indeed well-tuned, it should assign a higher probability to an ill-fitting but existing word which is still in distribution, than to a completely ill-fitting string that is out-of-distribution. In any case, BPE models catch up to character models if (and only if) provided with additional syntactic/semantic context information. While random context somewhat aids BPE models, a substantial gap remains between the largest BPE models and the character models, where performance remains excellent, even in the presence of implausible contexts.
@@
Figure 2 displays learning curves for syntactic agreement phenomena, lexical decision, and both surprisal conditions across the 19 saved checkpoints (complete curves reproduced in Appendix C). The first 10 checkpoints correspond to the first 10% of pretraining, the remaining 9 checkpoints represent 10% of training steps over the remaining 90% of pretraining. For character models, the high-frequency, low-frequency, and syntactic curves are clearly separated. On high-frequency data, word learning is rapid and follows power-law curves; the low-frequency scores improve a little later and at a lower rate, but with the same trajectory. Syntactic phenomena improve later, mostly in s-shaped curves (e.g., det.-noun agreement), or are not learned at all in small models (subj.-verb agreement). In contrast, the syntactic and lexical curves for BPE models form sheaves of s-shaped trajectories. There is no principal difference in learning dynamics between the syntactic and the word level, improvements [^2] We calculate Spearman-rank correlation between ordered accuracies for word-level tasks and BLiMP in Appendix D. occur simultaneously. This further confirms the results of our previous experiments: in BPE models, word learning is dependent on syntax learning, and words cannot be recognized reliably outside of plausible contexts. Additionally, these different levels of learning cannot be disentangled in BPE models, whereas in character models, syntactic learning follows word learning. [^3] Interestingly, the learning process is not finished when accuracy curves stabilize, cf. Appendix F.
5 Discussion
In contrast to previous work, this study set out to explore when (and if) LMs learn what valid words are, and not how LMs learn when words are validly used. How should these results now be interpreted in the light of language acquisition and the use of BabyLMs to model corresponding processes? In reality, words are not learned in isolation, but from usage. Yet, words have been widely shown to be represented as solid “standalone” units in the mental lexicon, and to be units that human learners acquire early on (cf. Waxman and Gelman, 2009, also Montag et al., 2018). Usage-based approaches have tested aspects of word learning independently from syntax (mostly in object naming tasks, cf. Tomasello and Todd, 1983) and relate it to pragmatic aspects of communication. For example, children only react to words they know and ignore similar-sounding words. They even struggle with learning words that are phonetically extremely similar (like our stimuli), and only later gain this capacity Stager and Werker (1997). Similarly, in production, the earliest words come in isolation, slowly emerge into pivot schemas and holophrases Tomasello (1992, 2003), and only then finally turn into complex sentences. Of course, syntax also aids in discovering aspects of “wordiness”, like SV(X)-sentences offering cues for agent-patient relationships. It would be very interesting to further disentangle these levels of word knowledge in a follow-up study, but our current study is not concerned with the “you shall know a word by the company it keeps”-level of word knowledge, but rather with the “what do valid words of the language look like”-level, which can be assessed via lexical decision. As such, we believe that the separated learning curves of our character-level models represent more human-like learning than the highly correlated curves found in the subword models, but in reality, an overlap between them is definitely expected.
Reasons for the tremendous performance differences between subword and character models remain open to further inquiry. One plausible explanation is that character models have much more context available to calculate meaningful sequence probabilities, as words are split into many more tokens. While this is true, it is also exactly the point that we are stressing here: it is hard to imagine that arbitrary subword units lead to human-like, plausible word-level representations (like in exemplar models of lexical storage, cf. Bybee, 2010), whereas character models might offer a more justified level of granularity (and, e.g., better fit reading times, cf. Oh et al., 2021). Our findings also align with results on LMs’ sensitivity to character-level perturbations Moradi and Samwald (2021); Zhu et al. (2024) and their inability to solve character-level tasks, like counting occurrences of the letter r in the word strawberry Zhang and He (2024); Shin and Kaneko (2024); Cosma et al. (2025).
6 Conclusion